OpenStudy (anonymous):

Eliminate the parameter. x = t + 4, y = t^2

OpenStudy (perl):

we can 'eliminate' the parameter t by algebra x -4 = t now substitute into y

OpenStudy (anonymous):

So it becomes x = -4 + 4 and y = -4^2? @perl

OpenStudy (anonymous):

The answers are: y = √x + x + 4 y = √x - 4 y = x^2 + 16 y = x^2 - 8x + 16 I chose y = x^2 + 16 as the answer. Is that right? @AnswerMyQuestions @perl @iGreen

OpenStudy (anonymous):

@brainbuzzed5646 @TheSmartOne @rose18920 any help?

OpenStudy (anonymous):

sorry @sunshine201 math is not my best subject.... i don't want to give you a wrong answer but if no one else answers you back you should just go with your own answer :)

OpenStudy (anonymous):

Okay :) thank you anyway.

OpenStudy (perl):

we can 'eliminate' the parameter t by algebra first solve for t x -4 = t now substitute into y = t^2 y = (x-4)^2

OpenStudy (perl):

now you can expand that

OpenStudy (anonymous):

So the answer would be y = x^2 - 8x + 16?

OpenStudy (anonymous):

@perl